OpenStudy (anonymous):

The mean weight of quarters in circulation must be estimated. How many quarters must we randomly select if we want to be 95% confident that the sample mean is within 0.025 g of the true population mean for all quarters? Based on results from a sample of quarters we can estimate the standard deviation as 0.068 g. All I really need is the margin of error (e) and I can go from there. But feel free to post the formula so I can be sure. Thanks in advance!

OpenStudy (anonymous):

::sniffle::

OpenStudy (kropot72):

\[\bar{x}+1.96\frac{\sigma}{\sqrt{n}}-(\bar{x}-1.96\frac{\sigma}{\sqrt{n}})=2 \times 0.025\] \[3.92\frac{\sigma}{\sqrt{n}}=0.05\ ..............(1)\] Now you can plug the value of sigma into equation (1) and solve for the sample size n.